OpenStudy (xixi743):

Differentiate the following with respect to x: f(x)= cos^2(x^2) - x - sin^3(x^3)

OpenStudy (xixi743):

\[\cos^2(x^2)-x-\sin^3(x^3)\] this is it if this is easier to look at.

OpenStudy (astrophysics):

Here set it up as such, so you can see clearly how to apply the differentiation rules \[f(x) = \left( \cos(x^2) \right)^2-x-\left( \sin(x^3) \right)^3\] should make your life easier now to differentiate.

OpenStudy (astrophysics):

Notice what the exponents really mean

OpenStudy (xixi743):

So I pull them in front? \[f(x)= 2(cosx^2)-x-3(sinx^3)^2\]

OpenStudy (astrophysics):

Yes, exactly we use the power rule, but you still have to use chain rule!

OpenStudy (xixi743):

oh shoot

OpenStudy (astrophysics):

So, \[f'(x) = 2\cos(x^2) \times (\cos(x^2))' -1-3(\sin(x^3))^2 \times (\sin(x^3))'\]

OpenStudy (xixi743):

what is the derivative of cosx^2? I'm not sure what happens with the x^2. I know cos becomes -sin...

OpenStudy (astrophysics):

You have to use chain rule again for those :P

OpenStudy (xixi743):

please tell me it becomes -2sinx

OpenStudy (xixi743):

This seems pretty long. Next step is probably to simplify it... \[f'(X) = 2(\cos^2)(-2sinx)-1-3(sinx^3)^2(3cosx^2)\]

OpenStudy (astrophysics):

\[f'(x) = \color{blue}{2\cos(x^2) \times - \sin(x^2) \times (x^2)'} -1-\color{red}{3\sin(x^3))^2 \times \cos(x^3) \times (x^3)'}\]

OpenStudy (xixi743):

oh god

OpenStudy (astrophysics):

I mean the derivatives shouldn't be too bad, just make sure you do it step by step, so you don't make mistakes.

OpenStudy (xixi743):

you don't pull the 2 in front of the sin I guess

OpenStudy (xixi743):

Lol this is a test question that I have to correct because I scored a C #oops

OpenStudy (astrophysics):

It should be -2xsin(x^2)

OpenStudy (astrophysics):

You're multiplying so it doesn't matter

OpenStudy (xixi743):

This is old material so i'm having to relearn some of it. Just finished studying MTV and Rolle's and L'Hospital and optimization.

OpenStudy (astrophysics):

Ah, good ole calc 1

OpenStudy (xixi743):

So was the previous equation I had written down correct?

OpenStudy (xixi743):

Well the class I'm taking now is AP Calc AB. I'm still in High school

OpenStudy (astrophysics):

Differentiation is very important in all of those haha, so it's crucial you understand it, it'll get simpler as you practice

OpenStudy (xixi743):

lol yeah I get pretty lost with all this. I didn't retain anything from precalculus so that didn't help

OpenStudy (astrophysics):

Not really :P you didn't take the chain rule the second time and you forgot to include the x^2 and x^3 terms in your terms

OpenStudy (astrophysics):

It's ok, I never took pre calc and jumped into calculus and got wrecked

OpenStudy (astrophysics):

Took me a while to get it right, so don't get too worried

OpenStudy (xixi743):

okay so now I have \[f'(x)= 2(cosx^2)(-sinx^2)(2x)-1-3(sinx^3)^2(cosx^3)(3x^2)\]

OpenStudy (astrophysics):

looks good!

OpenStudy (xixi743):

Now I must struggle through simplifying this if possible

OpenStudy (astrophysics):

Haha have fun, some of these may come in handy http://www.purplemath.com/modules/idents.htm

OpenStudy (xixi743):

Ah I wasn't sure what to combine besides the 2 and the 2x in the first part of the function

OpenStudy (xixi743):

Thanks, but I already have all of those memorized. I just struggle with applying them.

OpenStudy (astrophysics):

Well keep practicing then check with http://www.wolframalpha.com/

OpenStudy (xixi743):

That is a beautiful math tool. Thank you.

OpenStudy (astrophysics):

Yup does wonders :)